OpenStudy (anonymous):

Which of the following expressions is equal to (x+2)+(x-2)(2x+1)?

myininaya (myininaya):

(x+2)(1+(2x+1))=(x+1)(2x+2)=2(x+1)(x+1)=2(x+1)^2 are all the same expression

OpenStudy (anonymous):

so what would be the answer???

myininaya (myininaya):

i don't know what the following expressions

OpenStudy (anonymous):

oh.. sorry hold on

OpenStudy (anonymous):

2x2 -2x 2x2-4x 2x2 +x 4x2+2x

OpenStudy (anonymous):

the 2 are supposed to be squared

OpenStudy (anonymous):

*x2

myininaya (myininaya):

2(x^2+2x+1)=2x^2+4x+2 but i don't see this listed above

myininaya (myininaya):

oh wait a minute i can't read sometimes forget everything

OpenStudy (anonymous):

ok...

myininaya (myininaya):

ok it is 2x^2-2x

myininaya (myininaya):

(x+2)+(2x^2+x-4x-2)=x+2+(2x^2-3x-2)=2x^2-2x

OpenStudy (anonymous):

ok thankx

myininaya (myininaya):

The reason I said forget everything is because I thought I seen a (x+2) in each term sorry about that